OpenStudy (timaashorty):

Challenge Problem: Each Letter stands for a digit. Find out what number each letter stands for using the equation below. C + C + C + C + R = FF F * R= MF C * C = MR So C= ? R= ? F= ? M= ? (fourth grade challenge; trying to help my little sister)

OpenStudy (timaashorty):

anybody? @timo86m @shamil98 , @Zale101 @iambatman

OpenStudy (zale101):

well, is there a number at least given ? that fourth grade work seems challenging lol

OpenStudy (timaashorty):

Nope no numbers lol

OpenStudy (anonymous):

Are you sure no numbers were given o_O

OpenStudy (timaashorty):

nope no numbers, want a picture for proof ? lol

OpenStudy (anonymous):

yesh :3

OpenStudy (timaashorty):

you idiot you don't believe me. lmao gimme a sec.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I believe you

OpenStudy (anonymous):

This is one of those problems where the answer is right in front of you but you end up over complicating it.

OpenStudy (timaashorty):

yupp, what if they all equal 1 ?I mean if they do, it'll make sense

OpenStudy (anonymous):

I feel sick

OpenStudy (anonymous):

And if Fiz gets this answer i'm going to be pissed

OpenStudy (the_fizicx99):

Hmm have you tried assigning C, R F,and M a value? C + C + C + C = C^4 C^4 + R = F if F were 18, c was 2 and r was 2 then.. 2^4 + 2 = 18 but they would also have to match the other problems. Try finding a number that satisfy all of the problems >.<

OpenStudy (the_fizicx99):

;-; if fiz got it right I would also be shocked >.<

OpenStudy (anonymous):

Yeah I'm thinking you have to assign the digits yourself XD

OpenStudy (the_fizicx99):

o.0

OpenStudy (timaashorty):

LOl piss him off.

OpenStudy (the_fizicx99):

Waiiit c + c + c + c isn't c^4 >_<

OpenStudy (the_fizicx99):

that would onyl be c * c * c * c >_<

OpenStudy (the_fizicx99):

only*

OpenStudy (the_fizicx99):

Sheit I messed up on adding them ;-; im terra bad U_U you get the idea though

OpenStudy (timaashorty):

The proof. @fiz you're making sense but that is too much to solve for a 4th grader. I think it is a simple equation just very confusing.

OpenStudy (anonymous):

an acceptable answer is probably C = 1, R = 1, M = 1 and F = sqrt(5)

OpenStudy (timaashorty):

oh dang square roots, she haven't learned that, but I see what you're doing

OpenStudy (the_fizicx99):

Pfft, this is just adding >_< adding roots *cough* seems complicated for a 4th grade -looks away- lol but seriously just find some values... I think that FF is a double digit meaning like FF like 22 Or f*f which if it resulting in 22 divide by 2 and get 11. So each f = 11 And then you plug in 11 for the other f's and find a R value for example :D

OpenStudy (the_fizicx99):

@mathmale I'm not making sense honestly... I'm multi-tasking D:

OpenStudy (anonymous):

just reposting correct version second row: R = M third row: C^2 = R^2 = M^2 C = R = M first row: 5C = F^2 equally, 5R = F^2 and 5M = F^2 \[F = \sqrt{5C} = \sqrt{5M} = \sqrt{5R}\] where C, R and M are any number except 0 (because we divided by F to make our assumption in the second row; therefore F can't equal zero

OpenStudy (zale101):

do fourth graders learn about square roots?

OpenStudy (timaashorty):

Whoopwhoop got to give away two medals this time. See, if there was no M this would be much easier. D; But I'll take what you both did and figure it out with her.

OpenStudy (timaashorty):

nope I asked her, she has no idea what they are

OpenStudy (the_fizicx99):

k

OpenStudy (anonymous):

This is not 4th grade math o_O

OpenStudy (anonymous):

Maybe for an alien race

OpenStudy (timaashorty):

and batman seen the proof lol ?

OpenStudy (timaashorty):

YES it is. it's a challenge homework

OpenStudy (zale101):

i think her teacher assigned the class the numbers or something

OpenStudy (timaashorty):

Thanks Fizi (:

OpenStudy (anonymous):

oh wow. EACH LETTER STANDS FOR A DIGIT ignore my answer

OpenStudy (the_fizicx99):

*Cough* yw

OpenStudy (the_fizicx99):

LOL if her teacher gave her the numbers and she comes with a long equation high schooled based lmao her reaction X_X

OpenStudy (timaashorty):

@ zale ; nope, lol euler @tHe_FiZiCx99 haha quit coughing.

OpenStudy (anonymous):

Prob why she's teaching grade 4 math

OpenStudy (anonymous):

err he/she

OpenStudy (anonymous):

lol! yea this is impossible. especially or fourth grade

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

I repeat: This is one of those problems where the answer is right in front of you but you end up over complicating it.

OpenStudy (timaashorty):

I'll just put all equal 1, and tomorrow she'll ask the teacher, ill post what was the answer. lol

OpenStudy (the_fizicx99):

baman I would have seen it >_< like im just op :*

OpenStudy (anonymous):

I'm more OP, dem gadgets

OpenStudy (timaashorty):

Guys 1 won't work Dx

OpenStudy (timaashorty):

WTH.

OpenStudy (the_fizicx99):

1 is too easy >_<

OpenStudy (timaashorty):

and this question is too confusing. I thought maybe it could be one since this is 4th grade.

OpenStudy (anonymous):

I know how to do it

OpenStudy (timaashorty):

How??

OpenStudy (anonymous):

Brb

OpenStudy (timaashorty):

alright

OpenStudy (anonymous):

ill bbs and try it later

OpenStudy (timaashorty):

Lol tagg me if you ever got it.

OpenStudy (doc.brown):

Is it a word problem? C+C+C+C+R = Foreseer

OpenStudy (whpalmer4):

C=4, R = 6, F = 2, M = 1

OpenStudy (timaashorty):

Nope it just says FF

OpenStudy (timaashorty):

OMG you're are brilliant! At first I thought ( ff means f * f) I plugged it like this 4 + 4 + 4 + 4 + 6 = 2 x 2 0r 4 2 x 6 = 1x2 4 x 4 = 1x 6 But now I see 4 + 4 + 4 + 4 + 6 = 2 2 2 x 6 = 12 4 x 4 = 1 6 Haha Thank you so much (:

OpenStudy (anonymous):

whpalmer, how did you solve this?

OpenStudy (anonymous):

Oh Timaa lol, but yup @whpalmer4 is correct :P

OpenStudy (anonymous):

Took fiz's medal away and gave it to whp

OpenStudy (anonymous):

i suppose is trial and error then

OpenStudy (whpalmer4):

not quite, I'll write up my method, just a minute...

OpenStudy (whpalmer4):

So, starting with the C + C + C + C + R = F F or 4C + R = F*10 + F trying out the various values of C from 0 to 9, and then adding the possible values of R from 0 to 9 to that, and looking for a sum that has two identical digits, the possibilities are: C = 8 R = 1 F = 3 C = 5 R = 2 F = 2 C = 2 R = 3 F = 1 C = 7 R = 5 F = 3 C = 4 R = 6 F = 2 C = 1 R = 7 F = 1 C = 9 R = 8 F = 4 C = 6 R = 9 F = 3 Now let's discard all the cases where we have the same digit assigned to multiple letters: C = 8 R = 1 F = 3 C = 2 R = 3 F = 1 C = 7 R = 5 F = 3 C = 4 R = 6 F = 2 C = 9 R = 8 F = 4 C = 6 R = 9 F = 3 Our next constraint is C*C = M*10+R Going through our candidates: 8*8 = 64, but R = 1, not 4 2*2 =4, but R = 3, and we don't have a tens digit, 7*7 = 49, but R = 5, not 9 4*4 = 16, R = 6, good, and implies M = 1 9*9 = 81, R = 8, not 1 6*6 = 36, R = 9, not 6 So C = 4, R = 6, M = 1, and F = 2

OpenStudy (anonymous):

I still find it surprising they would assign this to a 9 year old lol.

OpenStudy (anonymous):

at first, i thought i was dealing with roman numerals haha...guess not... awesome solution dude

OpenStudy (whpalmer4):

Yeah, I haven't met anyone at my kid's school who could have done this at age 10 or 11, I think, much less at 9.

OpenStudy (anonymous):

you dont need to test all r's only 1,4,6,9,0

OpenStudy (whpalmer4):

solution is unique.

OpenStudy (anonymous):

and r equal 4 also implies in M=6

OpenStudy (timaashorty):

See, I kind of did what you did except I kept thinking FF = F*F, MR = M*R, & MF = M*F So I never really got the solution. Thanks for explaining it really well, I showed my sister this morning she got all excited when it was solved. lol & Batman me neither, but the teacher put it for her as a challenge/ bonus grade for fun I guess. It wasn't something they had to do.

OpenStudy (the_fizicx99):

I thought it was either f*f or a double meaning like 2 number >_<